# Demystifying Language Model Forgetting with Low-rank Example Associations

### v2

@@ Abstract
Large Language models (LLMs) suffer from forgetting of upstream data when fine-tuned. Despite efforts on mitigating forgetting, few have investigated whether, and how forgotten upstream examples are dependent on and associated with newly learned tasks. Insights on such associations enable efficient and targeted mitigation of forgetting. In this paper, we empirically analyze forgetting (measured in log-perplexity increase) that occurs in $N$ upstream examples of language modeling or instruction-tuning after fine-tuning LLMs on one of $M$ new tasks, visualized in $M\times N$ matrices. We demonstrate that the matrices display simple low-rank patterns, often well-approximated with multiplicative scalar effects of upstream examples and newly learned tasks. We also examine fine-grained associations with visualization and statistics. Leveraging the low-rank nature of the associations, we predict forgetting of upstream examples when fine-tuning on unseen tasks with matrix completion over the empirical associations. This enables fast identification of most forgotten examples without expensive inference on the entire upstream data. The approach, despite simplicity, outperforms prior approaches that learn semantic relationships of learned tasks and upstream examples with LMs for predicting forgetting. We demonstrate the practical utility of our analysis by showing statistically significantly reduced forgetting as we upweight predicted examples for replay at fine-tuning. Project page: https://inklab.usc.edu/lm-forgetting-prediction/
@@ 1 Introduction
There has been a growing need for long-term usability of LLMs. Although fine-tuning allows efficient and incremental adaptation of models, it risks catastrophic forgetting (McCloskey & Cohen, 1989; Goodfellow et al., 2014) of upstream knowledge learned in the pre-training or instruction-tuning phase, causing unintended prediction changes over known information. This is problematic for the stability of online deployed LLM systems, limiting the feasibility of continual fine-tuning in practice (Raffel, 2023; Shi et al., 2024).
@@
In this paper, we empirically study the associations between learned tasks and forgotten upstream examples of language modeling or instruction-tuning. We experiment with OLMo-1B, OLMo-7B, OLMo-7B-Instruct (Groeneveld et al., 2024) and MPT-7B (Team, 2023) models where upstream data is open-source. We perform statistics of forgetting (in log perplexity increase) over $N$ upstream examples, after fine-tuning the model on one of $M$ unseen instruction-tuning tasks, and represent the statistics in a $M\times N$ matrix. Afterwards, we visualize the matrices and fit the observations with statistical models to analyze the associations. Figure 1 illustrates the analysis setup.
@@ 3.2 What are More Fine-Grained Associations?
SVD reveals fine-grained associations. We present $R^{2}$ of fitting the assoication with $Z_{r}$ with progressively larger rank $r$ in Figure 3 (b). We notice that for OLMo-7B and OLMo-7B-Instruct (full-parameter fine-tuning), $R^{2}$ quickly increases to 0.69 and 0.78 with $Z_{3}$ (reconstruction with up to the third component). We visualize the matrices reconstructed from each component $s_{k}\bm{\alpha}_{k}\bm{\beta}_{k}^{T}$ in SVD of $Z$. Figure 4 provides an example of patterns captured by the $k$-th component. We demonstrate how the visualization extracts patterns of forgetting that is conditional on the learned tasks.
@@ 3.3 How do the Similarity Measures Interpret the Associations?
Textual similarity. We measure textual similarity $z^{\textrm{text}}_{ij}$ between learned tasks and forgotten examples with TF-IDF vectorized over each pair of learned tasks $T_{i}$ and upstream examples $x_{j}$.
@@
Table 1: Spearman’s correlation coefficients between various measures of similarity and the actual forgetting on upstream examples after fine-tuning on a task.
Similarity Metrics Spearman’s $\rho$
Textual $z^{\textrm{text}}$ 0.004
Gradient-Weight Difference $z^{\textrm{g-w}}$ 0.005
Gradient-Gradient $z^{\textrm{g-g}}$0.007
Forgetting correlates poorly with similarity measures of learned tasks and upstream examples. We evaluate Spearman’s correlation between the actual forgetting $z_{ij}$ and the various similarity measures $\{z^{\textrm{text}}_{ij},z^{\textrm{g-w}}_{ij},z^{\textrm{g-g}}_{ij}\}$ on OLMo-1B and summarize the results in Table 1. We notice that none of the similarity measures correlates with the actual forgetting, with a correlation $\rho<0.01$. An interpretation of the low correlation is that the model weights deviate from the initial weights after fine-tuning to a region where first-order approximation of forgetting does not hold. We further visualize the matrices of $z^{\textrm{g-w}}_{ij}$ and $z^{\textrm{g-g}}_{ij}$ in Fig. 8 in Appendix and provide a side-by-side comparison with the matrices of forgetting $Z$. The visualization displays distinct patterns among the three matrices. These results imply that although the association matrices $Z$ display a simple pattern, they are not well-interpreted with common similarity measures of learned tasks and forgotten examples.
@@ 4.1 Training and Evaluation of Forgetting Prediction
Our goal is to accurately predict forgetting $z_{ij}$ over upstream examples $x_{1..M}$ when the model is fine-tuned on an unseen task $T_{j}$ with a prediction model $g$, without running expensive LLM inference on all $x_{1..M}$. To evaluate this, we create training and test splits by partitioning the set of fine-tuning tasks (noted as $\mathcal{T}_{\textrm{train}}$ and $\mathcal{T}_{\textrm{test}}$) and the rows of the association matrices $Z$. We further control whether the $\mathcal{T}_{\textrm{train}}$ and $\mathcal{T}_{\textrm{test}}$ are from the same category of the tasks to test both in-domain and out-of-domain generalization ability of the prediction models. For OLMo-1B and 7B experiments, we use FLAN and the in-domain tasks and Tulu and Dolly as out-of-domain testing tasks. For OLMo-7B-Instruct experiments, we use MMLU and BBH as in-domain tasks and use TruthfulQA and Dolly as out-of-domain testing tasks. Details about the tasks included in the training, in-domain testing, and out-of-domain testing sets are discussed in Appendix B.
@@
Matrix completion approaches. We run matrix completion algorithms including additive linear, SVD, and k-nearest neighbors (KNN) models. The additive linear and the SVD models are introduced earlier in Sec. 3. Given the seed forgetting $z_{i}^{\mathcal{S}}$ of a task $T_{i}\in\mathcal{T}_{\textrm{test}}$, KNN finds tasks from $\mathcal{T}_{\textrm{train}}$ that have similar patterns of forgetting over the seed upstream examples $\mathcal{S}$. KNN computes an average of forgetting of top-k similar tasks from $\mathcal{T}_{\textrm{train}}$ weighted by their similarity as the prediction of forgetting caused by $T_{i}\in\mathcal{T}_{\textrm{test}}$ on the upstream examples $x_{1..M}$.
@@ 4.2 Mitigating Forgetting with Predicted Forgetting
Baselines of mitigating forgetting. We compare with diverse strategies of selecting upstream examples for sparse replay. As a main comparator, we examine whether weighted sampling with the predicted forgetting improves over random sampling of upstream examples (Random). We compare with Maximally Interfered Retrieval (MIR) (Aljundi et al., 2019a), a selection strategy sharing the similar notion of importance that forgotten examples should be selected for replay. The approach performs bi-level sampling by selecting the most forgotten examples from a small random subset of upstream data (set as $|\mathcal{S}|=30$). In addition, we apply strategies that consider different definitions of upstream example importance. We examine an approach based on perplexity thresholds (PPL) (Marion et al., 2023), which samples upstream data of which the perplexity is around the median of the distribution. For OLMo-1B, we also sample replayed examples proportional to the gradient inner products (Grad-Prod) we evaluated in Sec. 3.3, in a similar vein to coreset selection approaches that utilize gradient information (Park et al., 2023; Xia et al., 2024).
@@ 4.3 Results of Predicting and Mitigating Forgetting
Results of predicting example forgetting. Table 2 summarizes the error of predicting example forgetting over tasks from the in-domain and out-of-domain test splits. We see matrix completion approaches consistently outperform the learned similarity model in the prior work. Among the three matrix completion approaches, we notice that KNN models in general achieve the lowest prediction error. Given this finding, we primarily leverage predictions given by the KNN model to reduce forgetting during fine-tuning.
@@
Mitigating forgetting with the predicted forgetting. We leverage the forgetting predicted by the KNN model and upweight examples during replay following our procedure in Sec. 4.2. Figure 6 summarizes log perplexity after fine-tuning over the held-out (never replayed) upstream data as we apply different upstream example selection approaches. We visualize the relative comparisons of these approaches to no forgetting of upstream examples (log perplexity before fine-tuning). We notice that example selection based on gradient inner products (Grad) or perplexity threshold (PPL), mainly applied for identifying important training data for a task in prior works, does not show improvement in mitigating forgetting compared to replaying random examples. This implies that the notions of example importance in these works are different from how easily the examples are forgotten. We also notice that MIR does not improve over random sampling in our setup, likely because of the small size of the retrieval candidates relative to the upstream examples. We see upweighting examples with ground truth forgetting (GT) consistently reduces forgetting compared to random examples. Replaying with KNN-predicted forgetting also consistently reduces forgetting, achieving statistical significance in 4 out of 5 tested setups.
Limitations of replaying with forgetting prediction. We note that our current practice of replaying with predicted forgetting requires two runs of fine-tuning: a replay-free run of fine-tuning after which the seed forgetting will be evaluated, and another run while replaying examples with the predicted forgetting. This creates computational overhead equivalent to one extra run of fine-tuning, which is identical to the MIR baseline. Nevertheless, the practice is still efficient when the training set of fine-tuning is much smaller than the upstream data. Future works can predict forgetting on-the-fly during fine-tuning, mitigating the overhead of fine-tuning the model twice.
@@ Appendix A Dataset, Model, and LM Training Details
Learned new tasks and their categorization. We summarize the list and the categorization of newly learned tasks in Tables 3 and 4 in our experiments. We also include the number of training examples and forgetting caused by each task averaged over all upstream examples.
@@ Appendix B Details of Forgetting Prediction and Replay
Data Splits for Predicting Example Forgetting. We mark the tasks used as in-domain test splits for predicting example forgetting (Sec. 4) in Tables 3 and 4. The train-test split for the in-domain tasks is randomly generated.
Training and evaluation details. We use Surprise Library 1.1.3333https://github.com/NicolasHug/Surprise/tree/v1.1.3 for additive linear, SVD, and KNN prediction models. For SVD, we set the dimension of the learnable features as 5. We train the regression models for 100 epochs over the association matrices. The other hyperparameters are left as default.
@@
Appendix C Additional Results about Example Associations

### v3

@@ Abstract
Large Language models (LLMs) suffer from forgetting of upstream data when fine-tuned. Despite efforts on mitigating forgetting, few have investigated whether, and how forgotten upstream examples are dependent on and associated with newly learned tasks. Insights on such associations enable efficient and targeted mitigation of forgetting. In this paper, we empirically analyze forgetting (measured in log-perplexity increase) that occurs in $N$ upstream examples of language modeling or instruction-tuning after fine-tuning LLMs on one of $M$ new tasks, visualized in $M\times N$ matrices. We demonstrate that the matrices display simple low-rank patterns, often well-approximated with multiplicative scalar effects of upstream examples and newly learned tasks. We also examine fine-grained associations with visualization and statistics. Leveraging the low-rank nature of the associations, we predict forgetting of upstream examples when fine-tuning on unseen tasks with matrix completion over the empirical associations. This enables fast identification of most forgotten examples without expensive inference on the entire upstream data. The approach, despite simplicity, outperforms prior approaches that learn semantic relationships of learned tasks and upstream examples with LMs for predicting forgetting. We demonstrate the practical utility of our analysis by showing statistically significantly reduced forgetting as we upweight predicted examples for replay at fine-tuning.
@@ 1 Introduction
There has been growing need for promptly updating LLMs to mitigate harmful behaviors, update outdated knowledge, and expand the set of application domains. Although fine-tuning allows efficient and incremental updates of models, it risks catastrophic forgetting (McCloskey & Cohen, 1989; Goodfellow et al., 2014) of upstream knowledge learned in the pre-training or instruction-tuning phase, causing unintended prediction changes over known information. This is problematic for the stability of online deployed LLM systems, limiting the feasibility of continual fine-tuning in practice (Raffel, 2023; Shi et al., 2024).
@@
In this paper, we empirically study the associations between learned tasks and forgotten upstream examples of language modeling or instruction-tuning. We experiment with OLMo-1B, OLMo-7B, OLMo-7B-Instruct (Groeneveld et al., 2024) and MPT-7B (Team, 2023) models where upstream data is open-source. We analyze forgetting (in log perplexity increase) over $N$ upstream examples, after fine-tuning the model on one of $M$ unseen instruction-tuning tasks, and represent the results in a $M\times N$ matrix. Afterwards, we visualize the matrices and fit the observations with statistical models to analyze the associations. Figure 1 illustrates the analysis setup.
@@ 3.2 What are More Fine-Grained Associations?
SVD reveals fine-grained associations. We present $R^{2}$ of fitting the assoication with $Z_{r}$ with progressively larger rank $r$ in Figure 3 (b). We notice that for OLMo-7B and OLMo-7B-Instruct (full-parameter fine-tuning), $R^{2}$ quickly increases to 0.69 and 0.78 with $Z_{3}$ (reconstruction with up to the third component). We visualize the matrices reconstructed from each component $s_{k}\bm{\alpha}_{k}\bm{\beta}_{k}^{T}$ in SVD of $Z$. Figure 4 provides an example of patterns captured by the $k$-th component in OLMo-7B-Instruct. We demonstrate how the visualization extracts patterns of forgetting that is conditional on the learned tasks. In Appendix D, we further examine interpretable patterns in the decomposition of $Z$ in OLMo-1B and 7B.
@@ 3.3 How do the Similarity Measures Interpret the Associations?
Textual similarity. We measure textual cosine similarity $z^{\textrm{text}}_{ij}$ between learned tasks and forgotten examples with TF-IDF vectorized features over each pair of learned tasks $T_{i}$ and upstream examples $x_{j}$. We also measure text representation similarity with final layer representations of OLMo-1B.
@@
Table 1: Correlations between various measures of similarity and the actual forgetting on upstream examples after fine-tuning on a task.
Pearson $\rho$ Spearman $\rho$
Textual (TF-IDF) -0.049 -0.035
Textual (Representation) 0.021 0.017
Gradient-Weight Diff. $z^{\textrm{g-w}}$0.003 -0.009
Gradient-Gradient Diff. $z^{\textrm{g-g}}$ 0.061 0.052
Forgetting correlates poorly with similarity measures of learned tasks and upstream examples. We evaluate correlations between the actual forgetting $z_{ij}$ and the various similarity measures $\{z^{\textrm{text}}_{ij},z^{\textrm{g-w}}_{ij},z^{\textrm{g-g}}_{ij}\}$ on OLMo-1B and summarize the results in Table $1$The table is updated from the earlier preprint (https://arxiv.org/abs/2406.14026v2) due to an implementation issue.. We notice that none of the similarity measures correlates with the actual forgetting, with a correlation $|\rho|<0.1$. An interpretation of the low correlation is that the model weights deviate from the initial weights after fine-tuning to a region where first-order approximation of forgetting does not hold. We further visualize the matrices of $z^{\textrm{g-w}}_{ij}$ and $z^{\textrm{g-g}}_{ij}$ in Fig. 8 in Appendix and provide a side-by-side comparison with the matrices of forgetting $Z$. The visualization displays distinct patterns among the three matrices. These results imply that although the association matrices $Z$ display a simple pattern, they are not well-interpreted with common similarity measures of learned tasks and forgotten examples.
@@ 4.1 Training and Evaluation of Forgetting Prediction
Our goal is to accurately predict forgetting $z_{ij}$ over upstream examples $x_{1..M}$ when the model is fine-tuned on an unseen task $T_{j}$ with a prediction model $g$, without running expensive LLM inference on all $x_{1..M}$. To evaluate this, we create training and test splits by partitioning the set of fine-tuning tasks (noted as $\mathcal{T}_{\textrm{train}}$ and $\mathcal{T}_{\textrm{test}}$) and the rows of the association matrices $Z$. We further control whether the $\mathcal{T}_{\textrm{train}}$ and $\mathcal{T}_{\textrm{test}}$ are from the same category of the tasks to test both in-domain and out-of-domain generalization ability of the prediction models. For OLMo-1B and 7B experiments, we use FLAN and the in-domain tasks and Tulu and Dolly as out-of-domain testing tasks. For OLMo-7B-Instruct experiments, we use MMLU and BBH as in-domain tasks and use TruthfulQA and Dolly as out-of-domain testing tasks. Details about the tasks included in the training, in-domain testing, and out-of-domain testing sets are discussed in Tables 7 and 8 in Appendix B.
@@
Matrix completion approaches. We run matrix completion algorithms including additive linear, SVD  [^2] We improved the results of SVD in Table 2 after more hyperparameter tuning in this version., and k-nearest neighbors (KNN) models. The additive linear and the SVD models are introduced earlier in Sec. 3. Given the seed forgetting $z_{i}^{\mathcal{S}}$ of a task $T_{i}\in\mathcal{T}_{\textrm{test}}$, KNN finds tasks from $\mathcal{T}_{\textrm{train}}$ that have similar patterns of forgetting over the seed upstream examples $\mathcal{S}$. KNN computes an average of forgetting of top-k similar tasks from $\mathcal{T}_{\textrm{train}}$ weighted by their similarity as the prediction of forgetting caused by $T_{i}\in\mathcal{T}_{\textrm{test}}$ on the upstream examples $x_{1..M}$.
@@ 4.2 Mitigating Forgetting with Predicted Forgetting
As we have discussed in Sec. 4.1, predicting forgetting with matrix completion requires seed forgetting $z^{S}$ to be evaluated. We consider an offline and an online variant of the approach. The offline variant performs a replay-free run of fine-tuning on the task $T_{i}$, after which the seed forgetting will be evaluated. We then perform another run of fine-tuning while replaying examples with the predicted forgetting. This creates computational overhead equivalent to one extra run of fine-tuning, but is still efficient when the training set of fine-tuning is considerably smaller than the upstream data. The online variant instead replays random examples for first 10% of fine-tuning steps, after which it evaluates seed forgetting and determine examples to be replayed in the rest of 90% steps. Compared to the offline variant, this mitigates the extra overhead of fine-tuning by trading off the prediction accuracy of forgetting.
Baselines of mitigating forgetting. We compare with diverse strategies of selecting upstream examples for sparse replay. We primarily examine whether weighted sampling with predicted forgetting statistically significantly improves over random sampling of upstream examples (Random). We also compare with Maximally Interfered Retrieval (MIR) (Aljundi et al., 2019a), a selection strategy sharing the similar notion of importance that forgotten examples should be selected for replay. The approach performs bi-level sampling by selecting the most forgotten examples from a small random subset of upstream data (set as $|\mathcal{S}|=30$). In addition, we apply strategies that consider different definitions of upstream example importance. We examine an approach based on perplexity thresholds (PPL) (Marion et al., 2023), which samples upstream data of which the perplexity is around the median of the distribution. For OLMo-1B, we also sample replayed examples proportional to the gradient inner products (Grad-Prod) we evaluated in Sec. 3.3, in a similar vein to coreset selection approaches that utilize gradient information (Park et al., 2023; Xia et al., 2024). As a reference, we also experiment with upweighting upstream examples with ground truth forgetting $z_{ij}$, which, however, can face computational efficiency issues in practice.
@@ 4.3 Results of Predicting and Mitigating Forgetting
Results of predicting example forgetting. Table 2 summarizes the error of predicting example forgetting over tasks from the in-domain and out-of-domain test splits. We see matrix completion approaches consistently outperform the learned similarity model in the prior work. Among the three matrix completion approaches, we notice that SVD models in general achieve the lowest prediction error. Besides, KNN in general outperforms additive linear and learned similarity models while being highly computationally efficient. Given this finding, we primarily leverage predictions given by the KNN and SVD models to reduce forgetting during fine-tuning.
@@
Mitigating forgetting with the predicted forgetting. We leverage the online or offline predicted forgetting by the KNN and SVD models to reweight examples during replay following our procedure in Sec. 4.2. Figure 6 summarizes log perplexity after fine-tuning over the held-out (never replayed) upstream data as we apply different upstream example selection approaches. We visualize the relative comparisons of these approaches to no forgetting of upstream examples (log perplexity before fine-tuning). We notice that example selection based on gradient inner products (Grad) or perplexity threshold (PPL), mainly applied for identifying important training data for a task in prior works, does not show improvement in mitigating forgetting compared to replaying random examples. This implies that the notions of example importance in these works are different from how easily the examples are forgotten. We also notice that MIR does not improve over random sampling in our setup, likely because of the small size of the retrieval candidates relative to the upstream examples. Upweighting examples with ground truth forgetting (GT) consistently reduces forgetting compared to random examples. By utilizing predicted forgetting by offline SVD and KNN we statistically significantly reduce forgetting compared to random examples in 4 out of 5 tested setups. The nuanced differences between SVD and KNN in forgetting mitigation align with their in-domain prediction accuracy of forgetting in Table 2, where SVD outperforms KNN by most on OLMo-7B models. Utilizing online-predicted forgetting also statistically significantly improves over replaying random examples in 3 of the setups. The gaps between online and offline variants are closer on 7B models than 1B models.
Effects on downstream task performance measured with task-specific metrics. We evaluate fine-tuned OLMo-7B or OLMo-7B-Instruct on unseen LLM leaderboard tasks and present the results in Appendix C. We notice the performance stays stable or improves on some downstream tasks while degrades on others, indicating forgetting. Although we observe slightly improved performance of Offline-KNN to random or no replay on most forgotten tasks (e.g., Sciq on OLMo-7B and IFEval on OLMo-7B-Instruct), we do not see statistical significance. We leave effective algorithms to mitigate downstream task forgetting with predicted forgetting as future work.
Table 3: Computational cost of replay-based approaches as a summation of fine-tuning costs $FT(\cdot)$, inference costs over upstream examples $EV(\cdot)$, and matrix completion costs $MC$.
Method Cost
Random $FT(Y)$
Ground Truth $2FT(Y)+EV(N)$
Offline SVD, KNN $2FT(Y)+EV(S)+MC$
Online SVD, KNN $FT(Y)+EV(S)+MC$
MIR $2FT(Y)+Y\cdot EV(S)$
PPL,GradProd $FT(Y)$
Computational Efficiency. Table 3 summarizes the computation cost of the approaches as a function $FT(\cdot)$ of fine-tuning steps, a function $EV(\cdot)$ of upstream examples whose perplexity is evaluated, and the cost of matrix completion (MC) that is much smaller than LLM inference or training. We note the total number of upstream examples as $N$, the size of seed examples as $S$, and the number of fine-tuning steps as $Y$. As $S$ is much smaller than $M$, the majority of computational costs arise from fine-tuning $FT(Y)$ and $EV(N)$. Replaying with ground truth forgetting is the most costly, as it introduces an additional run of fine-tuning (after which forgetting will be evaluated) and inference over potentially very large-scale upstream data. As offline prediction of forgetting only mitigates the need for inference, the approach saves computations when the cost of fine-tuning $FT(Y)$ is notably smaller than $EV(N)$, i.e., over small fine-tuning datasets and massive upstream data. Online prediction of forgetting is always efficient, requiring only one run of fine-tuning and without the need to evaluate over massive upstream data.
@@ 6 Conclusions
Limitations. In our research process, we tried to control confounding factors such as model size, learning rate, or the number of training steps. However, we did not systematically study how the associations between learned tasks and forgotten upstream examples depend on these factors. For example, although we presented results of 1B and 7B models from the same family (OLMo) and observe differences in their association statistics (such as $R^{2}$ scores in Sec. 3.1), it is unclear how the associations will appear in smaller or larger models. Besides, our empirical findings raise questions about interpretability of the associations, i.e., why certain examples are more prone to forgetting while learning a new task, and whether there exist more human-interpretable patterns behind the associations. Finally, we limited our experiments to fine-tuning on a single task at a time. Predicting and mitigating forgetting in sequentially learned tasks is left as future work.
@@ Appendix A Dataset, Model, and LM Training Details
Learned new tasks and their categorization. We summarize the list and the categorization of newly learned tasks in Tables 7 and 8 in our experiments. We also include the number of training examples and forgetting caused by each task averaged over all upstream examples.
@@ Appendix B Details of Forgetting Prediction and Replay
Data Splits for Predicting Example Forgetting. We mark the tasks used as in-domain test splits for predicting example forgetting (Sec. 4) in Tables 7 and 8. The train-test split for the in-domain tasks is randomly generated.
Training and evaluation details. We use Surprise Library 1.1.3555https://github.com/NicolasHug/Surprise/tree/v1.1.3 for additive linear, SVD, and KNN prediction models. For SVD, we set the dimension of the learnable features as 5. We train the regression models for 100 epochs over the association matrices. The other hyperparameters are left as default.
@@
Appendix C Downstream Task Evaluation with Task-Specific Metrics
Table 4: Downstream task performance of OLMo-7B models before and after fine-tuning on dolly tasks.
ARC-Easy ARC-Challenge Boolq Hellaswag Openbookqa Piqa Sciq Winogrande
Metrics Acc-norm Acc-norm Acc Acc-norm Acc-norm Acc-norm Acc-norm Acc
Before FT 68.77 40.36 72.41 75.65 42.20 79.54 88.60 66.29
No Replay 67.34 42.28 74.82 76.89 44.65 80.05 84.09 67.89
Random 67.48 42.43 74.33 77.26 44.88 79.97 84.77 67.33
KNN-Offline 67.49 42.24 74.33 77.07 44.30 80.09 84.91 67.54
Table 5: Downstream task performance of OLMo-7B-Instruct models before and after fine-tuning on dolly tasks.
MMLU-Pro BBH IF-Eval MUSR GPQA
Metrics 5-shot Acc 3-shot Acc-norm 0-shot Inst 0-shot Acc-norm 0-shot Acc-norm
Before FT 18.2 20.27 39.93 20.65 14.34
No Replay 17.51 20.26 17.72 22.01 14.37
Random 17.45 20.02 18.19 21.90 14.41
KNN-Offline 17.45 20.19 18.28 21.90 14.42
We evaluate downstream task performance of LMs before and after fine-tuning on 8 tasks from Dolly with LM-Evaluation-Harness framework (Gao et al., 2024). For OLMo-7B models, we evaluate on the same set of downstream tasks in OLMo technical report (Groeneveld et al., 2024). For OLMo-7B-Instruct models, we evaluate on Open LLM Leaderboard tasks [^6] https://huggingface.co/docs/leaderboards/open_llm_leaderboard/about. For fine-tuned models, we compare no replay, replaying random examples, and replaying with forgetting predicted by offline KNN. Tables 4 and 5 summarize the results.
We notice that fine-tuning OLMo-7B on Dolly improves downstream performance on most of the downstream tasks. This aligns well with the purpose of fine-tuning a LM that is not instruction-tuned. Nevertheless, we notice performance degradation on two of the tasks, namely ARC-Easy and Sciq, which indicates forgetting. Although offline KNN achieves higher accuracy scores on these two tasks compared to no-replay (67.48 to 67.34, 84.91 to 84.09), we do not find the improvement statistically significant. For OLMo-7B-Instruct, fine-tuning on Dolly only improves performance on MUSR. The models clearly suffer from forgetting on the other tasks such as IFEval. Offline KNN achieves higher scores than random or no replay on IFEval (18.28 compared to 18.19 and 17.72), but we could not conclude about the significance of the improvement.
To summarize, we do not see clear performance improvement in downstream task performance (evaluated with task-specific metrics) by replaying random or chosen examples. We conjecture that replay-based approaches are not sufficient to mitigate forgetting on their own, and can be combined with other approaches such as careful learning rate scheduling or parameter regularization. We leave more effective algorithms to mitigate downstream task forgetting with predicted forgetting as future works.
Appendix D Towards Interpreting Fine-Grained Associations
Table 6: Semantic meaning of $k$-th component in the SVD of the association matrix $Z$. We identify top relevant learned tasks and upstream example domains to $k$-th component in the SVD of the association matrix $Z$.
OLMo-1B OLMo-7B
Learned Tasks More Forgotten Domain Learned Tasks More Forgotten Domain
flan/paws_wiki flan/squad_v2
$k=1$ flan/glue_mrpc None flan/fix_punct None
flan/story_cloze tulu/open_orca
flan/opinion_abstracts_idebate flan/mnli_matched
$k=2$ dolly/general_qa StackOverflow flan/mnli_mismatched None
flan/story_cloze flan/snli
flan/story_cloze flan/squad_v2
$k=3$ flan/fix_punct None flan/quac None
flan/true_case flan/fix_punct
math_dataset flan/rte
$k=4$ dolly/general_qa None flan/opinion_abstracts_idebate None
flan/opinion_abstracts_idebate flan/story_cloze